OpenStudy (bbb911):

This is Geometry: Find the area of a rectangle whose diagonal is 10 feet long and which makes a 55 degree angle with one of the sides.

OpenStudy (thewayiam983):

the rectangle is really two triangles, the diagonal being their hypotenuse thus, one side of the triangle is sin55=x/10, x =8.19 and the other is cos55=y/10, y=5.7 Thus, the area is xy=(8.19)(5.7)= 46.97

OpenStudy (bbb911):

okay heres the multiple choice for that q A. 81=(x+6)sq + (y-9)sq B. 7=(x+6)sq + (y-9)sq C. 28=(x-6)sq+ (y+9)sq D. 7=(x-6)sq + (y+9)sq

OpenStudy (thewayiam983):

None of those are area formulas. Are you sure those are the right options?

OpenStudy (bbb911):

wait omg!! im sorry :( i was lookin at the wrong question heres the real multiple choice A. 100ft sq B. 56 ft sq C. 48ft sq D. 47ft sq

OpenStudy (thewayiam983):

47 square feet, D is your answer

OpenStudy (bbb911):

thank you so much